OpenStudy (anonymous):

To grow grass in Mel Smith's yard, a landscaper estimates that he will need 55 pounds of grass seed. Each pound will cost Mel $1.83. The landscaper will also charge Mel 6 hours of labor at $12.50 per hour. If Mel pays early, he will be entitled to a cash discount of $3.50. Assuming Mel takes advantage of the discount, how much does he pay the landscaper?

OpenStudy (anonymous):

55*1.83=100.65 6*12.5=75 100.65+75=175.65 175.65-3.50=172.15

OpenStudy (anonymous):

that answer the end

OpenStudy (anonymous):

yes. 172.15

OpenStudy (anonymous):

Total the following: Six hundred sixty-eight and eight hundred one thousandths Twelve and forty-nine hundredths Three and four thousandths Fifty-one hundredths Three hundred ten and four tenths

OpenStudy (anonymous):

anyone know that answer